A population has an average u=50 and standard deviation of ó=2.3.what is the probability that a random sample is size n=8 will have a mean between 52 and 54?


1
Expert's answer
2021-04-28T08:45:48-0400

Solution:

"\mu=50,\sigma=2.3,n=8"

"P(52\le X\le 54)=P(X\le 54)-P(X\le 52)
\\=P(z\le \dfrac{54-50}{2.3/\sqrt{8}})-P(z\le \dfrac{52-50}{2.3/\sqrt{8}})"

"=P(z\le 4.91)-P(z\le 2.459)
=1-0.9931=0.0069"